# SciCoQA: Quality Assurance for Scientific Paper–Code Alignment

### 26 Mar 2026 v2

@@ 3.2 Data Analysis
Fig. 3 analyzes several dimensions of SciCoQA. While our real-world data is mostly from computer science (specifically from AI and its subdomains such as ML, CV and NLP), the synthetic data contains papers and code from Electrical Engineering and System Science, Physics, Statistics, Quantitative Biology, and Mathematics, making the data diverse across computational sciences.
We further analyze the discrepancy type (as defined in §3), by annotating each discrepancy in the real set as Difference, Paper Omission, or Code Omission. For the synthetic data, we provide type definitions in the prompt and generate the label along with the discrepancy. We find that $55\%$ of discrepancies in the real data are Differences, which is also the majority class in the synthetic data, accounting for $78\%$ of the data.
Beyond the discrepancy type, we analyze the affected component within the research pipeline. We introduce a taxonomy of six discrepancy categories: Algorithm: changes to step order, operations, or core logic; Model: architectural or initialization changes; Loss: alterations to loss definitions or terms; Evaluation: modifications to evaluation logic, metrics, or scripts; Data: dataset usage, preprocessing, augmentation, or filtering; and Training: changes to the learning process, schedule, or optimization. We apply this taxonomy to real-world data and the CS subset of synthetic data, excluding the other domains where these ML-specific concepts may not always apply cleanly. Similar to the discrepancy type annotation, we manually label the real instances and automatically generate labels for the synthetic ones. In the real data, discrepancies in the Algorithm and Loss dominate ($25\%$ and $24\%$) whereas the synthetic data has $26\%$ of discrepancies in the Algorithm, and Model discrepancies are the second-highest category, at $21\%$.
Finally, we visualize the number of tokens in the prompt. Papers have a median of 14,20014{,}200, codebases 39,27239{,}272, and the combined prompt 56,90356{,}903 tokens. Further, $73$ out of $276$ papers have more than $100k$ tokens combined with their codebase, making SciCoQA a challenging task to measure the models’ long-context abilities. We analyze the data by programming languages (§B.3) and the synthetic code (§B.2) further in the appendix.
@@ 4 Experiments
Given a paper and its code, we prompt the model to generate a list of discrepancies between the two (prompt in §F.6). We then parse the model output into individual discrepancies. The generation prompt contains the same instructions as those used to construct the ground-truth discrepancies. For implementation details, we refer to §A. We further ablate the discrepancy prediction experiment by providing only the code as input. This allows us to quantify the contribution of the paper, distinguishing between discrepancies that require cross-modal reasoning and those that can be inferred from the code (e.g., through comments or readmes), or the model’s parametric knowledge of the paper.
@@
We employ LLM-as-a-Judge (Zheng et al., 2023) to evaluate whether the predicted discrepancy matches the reference. Inspired by Wei et al. (2025a), we use a reasoning model, specifically GPT-OSS 20B (OpenAI, 2025b) (prompt in §F.7), because it is open-weight, enabling reproducibility, and its favorable speed-performance trade-off. For the real data where two ground truth descriptions exists (generated by Gemini 3.1 Pro and GPT-5), we consider matches against either as correct. To verify the evaluation setup, we annotate the predictions from $20$ discrepancies, where $10$ originate from GitHub and $10$ from reproducibility papers. To streamline our evaluation, we compute the top three most similar predicted discrepancies per model using EmbeddingGemma (Vera et al., 2025). We then manually assess whether each prediction matches the reference discrepancy. In total, this yields 1,0391{,}039 annotations, of which $143$ predictions match the reference. On this data, the LLM judge achieves an F1 score of $87.5\pm 1.1$, showing strong alignment with our annotations. [^3] We observe a slight variance in predictions despite a fixed seed in vLLM, we therefore report the mean and standard deviation over five runs. We further validate the setup by running Qwen3 32B as an alternative judge. We find Cohen’s Kappa between them at $\kappa=83.8$ and an F1 score of $91.9$. However, since the F1 score of Qwen3 32B with our human evaluation is at $70.0$, we retain GPT-OSS 20B as a judge. While the evaluation is a binary classification (reference and predicted discrepancies match or do not match), we refer to the performance as recall, as the predictions may contain unannotated, but valid, discrepancies, and we only evaluate whether the annotated ones are detected.
@@
We evaluate several state-of-the-art model families, including commercial and open models, as well as reasoning, instruction-tuned, and code-specific variants, specifically: GPT-5, Gemini, GPT-OSS, Qwen3, DeepSeek R1, Nemotron, and Mistral (exact model variants in Table 2).
@@ 5 Results
Fig. 4 reports the recall of the top-performing models (full results in Table 9). Overall, Gemini 3.1 Pro and GPT-5 Mini perform best, achieving a recall of $46.7\%$, leaving considerable room for improvement. On the synthetic data, GPT-5 detects $70.0\%$ discrepancies. Crucially, we observe a strong correlation ($r=0.94$) between recall on real and synthetic data (see §C.4), validating the synthetic subset as a reliable proxy for real-world discrepancies. While general scaling laws can be observed (i.e., models trained with more compute and data perform better), GPT-5 Mini and GPT-OSS 20B are an exception in our task, outperforming their larger variants slightly. We attribute this, on the one hand, to the self-preference bias of the judge for GPT-OSS, rating its own outputs better (Panickssery et al., 2024), and on the other hand, to the model’s verbosity. GPT-5 Mini and GPT-OSS 20B on average make $5.3$ and $5.8$ predictions per paper, while GPT-5 and GPT-OSS 120B make only $4.6$ and $5.0$ predictions, giving the smaller variants a higher chance to match the discrepancy at the expense of precision. Additionally, we find GPT-5 Codex to be inferior, despite it being a larger model than GPT-5 Mini (Codex is based on GPT-5). While Codex is generally superior in code generation, for SciCoQA code and natural language understanding are both crucial, and we conjecture that the general instruction-following and reasoning abilities of GPT-5 and GPT-5 Mini are more helpful than specialized coding knowledge.
Origin & Type
@@
Analyzing detection rates reveals a clear hierarchy: Code Omissions are the easiest to detect, followed by Differences, while Paper Omissions are the most challenging. This order also explains the performance gap between data sources. Discrepancies from GitHub are easier to detect because they primarily consist of Differences ($71.4\%$) and Code Omissions ($19.0\%$). In contrast, discrepancies from reproducibility papers are more challenging as they are often dominated by Paper Omissions ($50\%$). This likely stems from the asymmetry between modalities: the paper acts as a specification, where all described components must exist in the code. Code Omissions and Differences benefit from this explicit grounding. Conversely, the code contains many implementation details not required in the paper; thus, detecting Paper Omissions requires the harder task of distinguishing deviations from permissible engineering artifacts without a textual reference.
Context Length
@@
We further analyze the performance by the number of input tokens.444§C.3 provides an analysis by relevant code file position. We split the dataset by prompt length into five approximately equally sized buckets. We find a consistent pattern where the longer the input, the lower the performance becomes, as similarly observed in various studies (Hsieh et al., 2024; Zhang et al., 2024; Levy et al., 2024). This sensitivity to input length also suggests an explanation for the performance between domains in the synthetic data: models struggle more with CS papers, which is likely driven by the larger repository size of CS vs non-CS repositories (median: $53k$ vs $31k$ tokens). This increased volume creates a larger search space, making the task more challenging. While many long context evaluations rely on artificially constructed “needles” that need to be retrieved from a long context, i.e., the haystack (Kamradt, 2023), SciCoQA can offer a natural evaluation setup similar to Liu et al. (2024a) for long-context evaluation of LLMs. [^5] The synthetic data generation can be used to inject discrepancies at controlled places in the codebase.
@@
Lastly, we split the data by publication year. Among the top models, the one with the most recent knowledge cutoff is the Gemini family, which was trained using data up to January 2025 (other models have cutoffs in 2024). Therefore, the 2025 split of our data can be considered not part of the pre-training data. Most models (except Gemini 2.5 Flash) perform worst on the most recent data. This demonstrates that models benefit from having the paper and code included in their pre-training data, aligning with previous findings on data contamination (OpenAI, 2023; Zhou et al., 2023). This further emphasizes the importance of our synthetic data generation pipeline, which enables us to continuously update our dataset with uncontaminated data that is not part of the pre-training data of future models.
Open-Models
Beyond the top-performing models, we find the other open-weight models (Qwen3, Nemotron, DeepSeek R1, Mistral) severely limited. Nemotron 49B and Qwen3 Coder 30B perform best, but only achieve an average recall on the combined data of $23.9\%$ and $23.5\%$, respectively.
In summary, SciCoQA provides a significant challenge for state-of-the-art LLMs. We find the recall of all models to be limited; the best-performing models can only detect $46.7\%$ of all discrepancies in the real-world data. Consequently, these systems cannot yet be relied upon to ensure the faithfulness of scientific publications with their codebases.
@@ 5.2 Validation of Unlabeled Discrepancies
To better understand the precision of the models, we analyze GPT-5’s, Gemini 2.5 Pro’s, and GPT-OSS 20B’s generations. We select predictions of $20$ papers from NLP and CV that were not matched to a ground truth, yielding a total of $224$ evaluations (results in Table 3). We select these papers to leverage our own domain expertise, as validating discrepancies is challenging and requires a detailed understanding of the paper and its implementation.
Overall, we find that Gemini 2.5 Pro does not make significant errors, achieving a precision of $94.1\%$. Its three errors were due to preprocessing (an omitted term in a formula during OCR), an ambiguity within the paper, a nuance even noted in the model’s generation and one error due to the incorrect recall of another paper’s formula. GPT-5 follows ($85.7\%$), primarily failing due to paper ambiguities or misinterpretations, or incorrect assumptions about library functions. Finally, GPT-OSS 20B struggles most ($68.0\%$), frequently misunderstanding the code logic (e.g., it fails to align variables that are named differently in the paper and code or misinterprets conditional execution paths).
Taking inspiration from information retrieval benchmarks where incomplete relevant judgments are common (Buckley and Voorhees, 2004), we construct a pooled ground truth from the $20$ papers. We aggregate the discrepancies originally annotated in SciCoQA with the verified ones identified by the models (excluding minor ones, which are mostly discrepancies related to mismatching settings that can typically be resolved via configuration or the CLI). This results in a comprehensive set of $129$ distinct paper-code discrepancies (results in Table 1). We find that the number of discrepancies uniquely identified by the models is high (see also Table 3), and recall remains low ($41$-$56\%$), confirming our main experiment’s findings that model recall is limited. GPT-5 achieves the best precision-recall tradeoff, obtaining an F1 score of $64.7\%$. We evaluate all $22$ models against this pooled ground truth in §C.6, confirming that recall remains the primary bottleneck.
@@
Table 1: Performance of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B on $129$ pooled discrepancies from the annotation of unlabeled predictions and discrepancies from SciCoQA from $20$ NLP and CV papers.
@@ 6 Conclusion
We introduced SciCoQA, a dataset of $635$ discrepancies designed to evaluate the alignment between a scientific paper and its code. Our analysis reveals a critical gap: while models like GPT-5 and Gemini demonstrate high precision, they suffer from insufficient recall, detecting only $46.7\%$ of real-world discrepancies. We find that models particularly struggle with paper omissions, where code logic is absent from the paper text, and fail to maintain performance on recent, uncontaminated publications. Consequently, while current LLMs show promise as assistants, they cannot yet serve as autonomous arbiters of scientific validity. As research scales toward “AI Scientists,” SciCoQA provides the essential ground truth to ensure these agents remain faithful to the scientific method. To further advance this field, we encourage the broader adoption of reproducibility tracks to systematically capture these discrepancies and expand future data collection.
@@ 7 Limitations
Our real-world data is predominantly skewed towards Computer Science and Artificial Intelligence. While we mitigate this by including synthetic discrepancies from Physics, Quantitative Biology, and Electrical Engineering, the distribution of errors in these fields may differ from our synthetic approximations. Consequently, model performance on non-CS domains should be interpreted with this synthetic nature in mind. We applied our real-world pipeline also to non-CS domains, searching GitHub repositories for URLs linking to PubMed, bioRxiv, medRxiv, but found very few with associated issues reporting discrepancies. Further, to our knowledge, no domain outside CS/AI systematically produces reproducibility papers with paper-code discrepancies. For this reason, we expanded to other domains via synthetic data, which we believe provides more value than excluding these domains entirely and we include a call for more reproducibility efforts in the conclusion.
@@
Our synthetic discrepancies are generated by GPT-5, which is also evaluated on the benchmark. We find that GPT-5 achieves disproportionately higher recall on synthetic data. This suggests that GPT-5 may benefit from its involvement in the data generation on this subset. However, the strong correlation between real and synthetic recall across all $22$ models ($r=0.94$, Fig. 9) confirms that relative model comparisons remain valid. Notably, excluding the GPT-5 family from the correlation analysis increases the correlation to $r=0.98$, indicating that GPT-5’s synthetic advantage is an outlier that slightly weakens the overall trend rather than inflating it. To mitigate this limitation, absolute model performance should be assessed on the real subset, while the synthetic subset reliably supports relative model comparison and allows for targeted generation of new, uncontaminated data.

### 22 Apr 2026 v3

@@ 3.2 Data Analysis
Domain
Fig. 3 analyzes several dimensions of SciCoQA. While our real-world data is mostly from computer science (specifically from AI and its subdomains such as ML, CV, and NLP), the synthetic data contains papers and code from Electrical Engineering and Systems Science, Physics, Statistics, Quantitative Biology, and Mathematics, making the data diverse across computational sciences.
Taxonomy
We further analyze the discrepancy type (as defined in §3) by annotating each discrepancy in the real set as Difference, Paper Omission, or Code Omission. For the synthetic data, we provide type definitions in the prompt and generate the label along with the discrepancy. We find that $55\%$ of discrepancies in the real data are Differences, which is also the majority class in the synthetic data, accounting for $78\%$ of the data. Beyond the discrepancy type, we analyze the affected component within the research pipeline. We introduce a taxonomy of six discrepancy categories, namely: Algorithm, Model, Loss, Evaluation, Data, and Training. Table 1 provides an overview of our proposed taxonomy. We apply this taxonomy to real-world data and the CS subset of synthetic data, excluding the other domains where these ML-specific concepts may not always apply cleanly. Similar to the discrepancy type annotation, we manually label the real instances and automatically generate labels for the synthetic ones. In the real data, discrepancies in the Algorithm and Loss categories dominate ($25\%$ and $24\%$), whereas in the synthetic data Algorithm is the largest category ($26\%$) and Model is the second-largest ($21\%$).
Label Definition
Type Difference The code implements logic distinct from the paper’s description
Paper Omission The code includes critical components not described in the paper
Code Omission A step described in the paper is absent from the code
Category Algorithm Changes to step order, operations, or core logic
Model Architectural or weight initialization deviations
Loss Alterations to loss definitions or terms
Evaluation Modifications to evaluation logic, metrics, or scripts
Data Dataset usage, preprocessing, augmentation, or filtering
Training Changes to the learning process, schedule, or optimization
Table 1: The SciCoQA discrepancy taxonomy. Types describe the structural nature of the mismatch; Categories describe which component of the research pipeline is affected.
Input Length
Finally, we visualize the token count distributions of the papers and codebases. The medians are 14,20014{,}200 tokens for papers, 39,27239{,}272 for codebases, and 56,90356{,}903 when concatenated. Further, $73$ out of $276$ papers have more than $100k$ tokens combined with their codebase, making SciCoQA a challenging task for measuring the models’ long-context abilities. In the appendix, we analyze the dataset by programming languages and the synthetic code diffs (§B.3 and §B.2).
@@ 4 Experiments
Given a paper and its code, we prompt the model to generate a list of discrepancies between the two (prompt in §F.6). We then parse the model output into individual discrepancies. The generation prompt contains the same instructions as those used to construct the ground-truth discrepancies. For implementation details, we refer to §A. We further ablate the discrepancy prediction experiment by providing only the code as input. This allows us to quantify the contribution of the paper, distinguishing between discrepancies that require cross-modal reasoning and those that can be inferred from the code (e.g., through comments or readmes) or from the model’s parametric knowledge of the paper.
@@
We employ LLM-as-a-Judge (Zheng et al., 2023) to evaluate whether the predicted discrepancy matches the reference. Inspired by Wei et al. (2025a), we use a reasoning model, specifically GPT-OSS 20B (OpenAI, 2025) (prompt in §F.7), because it is open-weight, enabling reproducibility, and offers a favorable speed-performance trade-off. For the real data where two ground truth descriptions exist (generated by Gemini 3.1 Pro and GPT-5), we consider matches against either as correct. To verify the evaluation setup, we annotate the predictions from $20$ discrepancies, where $10$ originate from GitHub and $10$ from reproducibility papers. To streamline our evaluation, we compute the top three most similar predicted discrepancies per model using EmbeddingGemma (Vera et al., 2025). We then manually assess whether each prediction matches the reference discrepancy. In total, this yields 1,0391{,}039 annotations, of which $143$ predictions match the reference. On this data, the LLM judge achieves an F1 score of $87.5\%\pm 1.1$, showing strong alignment with our annotations. [^4] We observe a slight variance in predictions despite a fixed seed in vLLM; we therefore report the mean and standard deviation over five runs. We further validate the setup by running Qwen3 32B as an alternative judge. We find a Cohen’s Kappa of $\kappa=0.838$ between them and an F1 score of $91.9\%$. However, since the F1 score of Qwen3 32B against our human evaluation is only $70.0\%$, we retain GPT-OSS 20B as a judge. While the evaluation is a binary classification (reference and predicted discrepancies match or do not match), we refer to the performance as recall, as the predictions may contain unannotated, but valid, discrepancies, and we only evaluate whether the annotated ones are detected.
@@
We evaluate several state-of-the-art model families, including commercial and open-weight models, as well as reasoning, instruction-tuned, and code-specific variants, specifically: GPT-5, Gemini, GPT-OSS, Qwen3, DeepSeek R1, Nemotron, and Mistral (exact model variants in Table 3).
@@ 5 Results
Fig. 4 reports the recall of the top-performing models (full results in Table 10). Overall, Gemini 3.1 Pro and GPT-5 Mini perform best, achieving a recall of $46.7\%$, leaving considerable room for improvement. On the synthetic data, GPT-5 detects $70.0\%$ of discrepancies. Crucially, we observe a strong correlation ($r=0.94$) between recall on real and synthetic data, validating the synthetic subset as a reliable proxy for ranking models (see §C.4). While general scaling laws can be observed (i.e., models trained with more compute and data perform better), GPT-5 Mini and GPT-OSS 20B are exceptions in our task, slightly outperforming their larger variants. We attribute this partly to the self-preference bias of the judge for GPT-OSS, rating its own outputs better (Panickssery et al., 2024), and partly to the model’s verbosity. GPT-5 Mini and GPT-OSS 20B on average make $5.3$ and $5.8$ predictions per paper, while GPT-5 and GPT-OSS 120B make only $4.6$ and $5.0$ predictions, giving the smaller variants a higher chance to match the discrepancy at the expense of precision. Additionally, we find GPT-5 Codex to be inferior, despite being a larger model than GPT-5 Mini (Codex is based on GPT-5). While Codex is generally superior in code generation, for SciCoQA, code and natural language understanding are both crucial, and we conjecture that the general instruction-following and reasoning abilities of GPT-5 and GPT-5 Mini are more helpful than specialized coding knowledge.
Origin and Type
@@
Analyzing detection rates reveals a clear hierarchy: Code Omissions are the easiest to detect, followed by Differences, while Paper Omissions are the most challenging. This order also explains the performance gap between data sources. Discrepancies from GitHub are easier to detect because they primarily consist of Differences ($71.4\%$) and Code Omissions ($19.0\%$). In contrast, discrepancies from reproducibility papers are more challenging as they are often dominated by Paper Omissions ($50\%$). The challenging nature of Paper Omissions likely stems from the asymmetry between modalities: the paper acts as a specification, where all described components must exist in the code. Code Omissions and Differences benefit from this explicit grounding. Conversely, the code contains many implementation details not required in the paper; thus, detecting Paper Omissions requires the harder task of distinguishing deviations from permissible engineering artifacts without a reference in the paper.
Input Length
@@
We further analyze the performance by the number of input tokens.555§C.3 conducts a further performance analysis by the position of the relevant code file in the prompt. We split the dataset by prompt length into five approximately equally sized bins. We find a consistent pattern where the longer the input, the lower the performance becomes, consistent with findings in prior work (Zhang et al., 2024; Levy et al., 2024). This sensitivity to input length also suggests an explanation for the performance gap between domains in the synthetic data: models struggle more with CS papers, which is likely driven by the larger repository size of CS vs. non-CS repositories (median: $53k$ vs. $31k$ tokens). This increased volume creates a larger search space, making the task more challenging. Unlike long-context evaluations that primarily test retrieval of isolated facts from extended contexts (Kamradt, 2023; Hsieh et al., 2024), SciCoQA requires cross-document reasoning over real scientific papers and multi-file codebases to identify semantic mismatches, making it a natural testbed for long-context abilities (similar to the question answering task of Liu et al. (2024a)). [^6] The synthetic data generation can further be used to inject discrepancies at controlled positions in the codebase, enabling targeted analysis of context-length effects.
@@
Lastly, we split the data by publication year. Among the top models, the one with the most recent knowledge cutoff is the Gemini family, which was trained using data up to January 2025 (other models have cutoffs in 2024). Therefore, the 2025 split of our data is largely outside the pre-training data, although Gemini models have partial coverage. Most models (except Gemini 2.5 Flash) perform worst on the most recent data. This suggests that models benefit from the inclusion of specific papers and codebases in their pre-training sets, lending support to concerns regarding data contamination discussed in the literature (OpenAI, 2023; Zhou et al., 2023; Sainz et al., 2023). This further emphasizes the importance of our synthetic data generation pipeline, which enables us to continuously update our dataset with uncontaminated data that is not part of the pre-training data of future models.
Open-Weight Models
Beyond the top-performing models, we find the other open-weight models (Qwen3, Nemotron, DeepSeek R1, Mistral) severely limited. Nemotron 49B and Qwen3 30B Coder perform best, but only achieve an average recall on the combined data of $23.9\%$ and $23.5\%$, respectively.
In summary, SciCoQA provides a significant challenge for state-of-the-art LLMs. We find the recall of all models to be limited; the best-performing models can only detect $46.7\%$ of all discrepancies in the real-world data. Consequently, these systems cannot yet be relied upon to ensure the faithfulness of scientific publications to their codebases.
5.1 Code-Only Ablation
@@ 5.2 Validation of Unlabeled Discrepancies
To better understand the precision of the models, we analyze GPT-5’s, Gemini 2.5 Pro’s, and GPT-OSS 20B’s generations. We select predictions from $20$ papers in NLP and CV that were not matched to a ground truth, yielding a total of $224$ evaluations. We select these papers to leverage our own domain expertise, as validating discrepancies is challenging and requires a detailed understanding of the paper and its implementation. We report results of this precision analysis in Table 4.
Overall, we find that Gemini 2.5 Pro does not make significant errors, achieving a precision of $94.1\%$. Its three errors were due to preprocessing (an omitted term in a formula during OCR), an ambiguity within the paper (a nuance even noted in the model’s generation), and the incorrect recall of another paper’s formula. GPT-5 follows ($85.7\%$), primarily failing due to paper ambiguities or misinterpretations, or incorrect assumptions about library functions. Finally, GPT-OSS 20B struggles most ($67.0\%$), frequently misunderstanding the code logic (e.g., failing to align variables named differently in the paper and code or misinterpreting conditional execution paths).
[…]
@@
Table 2: Performance of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B on $129$ pooled discrepancies from the annotation of unlabeled predictions and discrepancies from SciCoQA from 2020 NLP and CV papers.
@@ 6 Conclusion
We introduced SciCoQA, a dataset of $635$ discrepancies designed to evaluate the alignment between a scientific paper and its code. Our analysis reveals a critical gap: while models like GPT-5 and Gemini demonstrate high precision, they suffer from insufficient recall, detecting only $46.7\%$ of real-world discrepancies. For a verification use case, low recall is a particularly consequential failure mode: missed discrepancies silently provide false assurance, while false positives can be filtered by a human reviewer. We further find that models struggle with Paper Omissions, where code logic is absent from the paper text, and fail to maintain performance on recent, uncontaminated publications. Consequently, while current LLMs show promise as assistants, they cannot yet serve as autonomous verifiers of paper-code faithfulness. As research scales toward “AI Scientists,” SciCoQA provides the essential ground truth to ensure these agents remain faithful to the scientific method. Beyond this primary use case, SciCoQA can serve multiple roles: as a natural long-context reasoning benchmark complementing synthetic retrieval probes, as a testbed for cross-modal claim verification between text and code, and as an evaluation suite for code comprehension in scientific domains. We release the dataset, model generations, evaluation pipeline, and synthetic generation code to support these applications. To further advance this field, we encourage the broader adoption of reproducibility tracks to systematically capture these discrepancies and expand future data collection.
@@ 7 Limitations
Our real-world data is predominantly skewed towards Computer Science and Artificial Intelligence. Applying our real-world pipeline to non-CS domains, searching GitHub repositories for URLs linking to PubMed, bioRxiv, and medRxiv, yielded very few issues reporting paper-code discrepancies, and we did not identify comparable reproducibility tracks outside of CS with paper-code discrepancy reports that we could use. The infrastructure for surfacing paper-code discrepancies thus appears to be largely absent outside CS/AI. Our synthetic pipeline is a short-term response, which we believe provides more value than excluding these domains entirely. Long-term, we call for broader adoption of reproducibility tracks, journals, and projects, such as ReScience C (Rougier et al., 2017) and the new Replication Research journal (Röseler et al., 2025). Consequently, the distribution of errors outside CS/AI may differ from our synthetic approximations, and performance on non-CS domains should be interpreted with this synthetic nature in mind.
@@
Our synthetic discrepancies are generated by GPT-5, which is also evaluated on the benchmark. We find that GPT-5 achieves disproportionately higher recall on synthetic data. This suggests that, on this subset, GPT-5 may benefit from its involvement in the data generation. However, the strong correlation between real and synthetic recall across all $22$ models ($r=0.94$, Fig. 9) confirms that relative model comparisons remain valid. Notably, excluding the GPT-5 family from the correlation analysis increases the correlation to $r=0.98$, confirming that the overall trend is not driven by GPT-5’s generation advantage; rather, the GPT-5 family sits slightly off the shared trend line that the other models follow closely. To mitigate this limitation, absolute model performance should be assessed on the real subset, while the synthetic subset reliably supports relative model comparison and allows for targeted generation of new, uncontaminated data.